if three dice thrown simultaneously ,then probability of getting a score of 5 is ;


 


(a)   5/216


(b)   1/6


(c)    1/36


 

Grade:11

2 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Piyush,

Ans:- The sample spaces are

(1,1,3)..........thrice

(1,2,2)...........thrice

Total no of probable cases are=216

Hence the prob =6/216=1/6(ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Piyush !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR

swetha viriyala
13 years ago

answer is 1/36 since only 6 chances out of 216 are possible for getting a sum 5